What is the correct ranking of the magnitude of the electric field, from largest to smallest, due to the single point charges at the locations shown? 2q 2q Eb > Ea Ec > Ed Ec > Ea>Eb >Ed o Ea = Ec > Eb > Ed Ed > Ea>Eb > Ec Ec > Ea>Ed >Eb
Answer #1

\text{Electric field due to a point charge is given as} E=\frac{1}{4\pi\epsilon_\circ}\frac{q}{r^2}

With the above formula we can see

that

Eb>Ea>Ec>Ed

the figure is not showing d but the answer is correct

first option